# Bar formation and evolution in the cosmological context: Inputs from the Auriga simulations

### v2

@@ 5.3 The multi-parameter space of bar formation
We attempt to summarise our findings in Fig. 17, where we show the distribution of barred and unbarred galaxies in different combinations of the parameter space we have explored in the previous sections. We find that, as discussed, the main parameters that separate out the barred-unbarred population are mainly the baryon-dominance $f_{\rm bd}$, the fraction of accreted stars within the inner region (bulge) of the galaxy $f_{\rm acc}$ and the Toomre $Q$ parameter of the inner galaxy $Q_{r_{50}}$, at the time of bar formation, although we note that there is significant overlap between in which galaxy will be barred or unbarred in either one of these properties separately. However, for some of properties, for example the gas fraction $f_{\rm gas}$ or the dark matter halo mass $M_{\rm DM}$, the barred and unbarred samples are completely overlapping. In future, this multi-dimensional parameter space can be used to build a semi-empirical model for bar formation within the cosmological context.
@@ 6 Summary & Conclusions
  • The bar formation time correlates with the time at which the galaxy becomes baryon-dominated, as well as with the baryon-dominance of the galaxy at $z=0$, but does not correlate with the host galaxy’s stellar mass $M_{\rm\star}$ (e.g. Fig. 5 and Fig 6).
@@
The fact that barred galaxies tend to be more baryon-dominated has important consequences for the loci of barred and unbarred galaxies in the Tully-Fisher (TF) relation (see Fig.  16). Namely, barred galaxies are offset from unbarred galaxies in the TF relation (i.e. for the same $V_{\rm flat}$, barred galaxies have a larger $M_{\star}$ or $M_{\rm baryonic}$). Confirming this offset for barred galaxies in the observed TF relation would have important implications for the existence of dark matter, since such an offset cannot be readily explained in a modified gravity (e.g. MOND) framework.
@@ Appendix C The Baryonic Tully Fisher Relation
Figure 21: The Baryonic Tully Fisher relation for galaxies in Auriga. Left: The grey dot-dashed line shows the relation given by fitting all galaxies together, while the shaded region indicates the intrinsic scatter $\sigma_{y}$. Right: The red (blue) dashed line shows the relation when fitting only the barred (unbarred) galaxies. The shaded regions indicate the uncertainties from boostrapping (see text). The scatter is slightly reduced for the BTFR as compared to the STFR and the barred and unbarred BTFR relations are less separated.

### v3

@@ 5.3 The multi-parameter space of bar formation
We attempt to summarise our findings in Fig. 18, where we show the distribution of barred and unbarred galaxies in different combinations of the parameter space we have explored in the previous sections. We find that, as discussed, the main parameters that separate out the barred-unbarred population are mainly the baryon-dominance $f_{\rm bd}$, the fraction of accreted stars within the inner region (bulge) of the galaxy $f_{\rm acc}$ and the Toomre $Q$ parameter of the inner galaxy $Q_{r_{50}}$, although we note that there is significant overlap between in which galaxy will be barred or unbarred in either one of these properties separately. However, for some of properties, for example the gas fraction $f_{\rm gas}$ or the dark matter halo mass $M_{\rm DM}$, the barred and unbarred samples are completely overlapping. In future, this multi-dimensional parameter space could be used to build a semi-empirical model for bar formation within the cosmological context.
@@ 6 Summary & Conclusions
  • The bar formation time, $t_{\rm bf}$, correlates with: i) the time at which the galaxy becomes baryon-dominated, ii) the time at which the galaxy assembles 50% of its stellar mass, as well as with iii) the baryon-dominance of the galaxy at $z=0$; it does not correlate with the host galaxy’s stellar mass $M_{\rm\star}$ at $z=0$ (e.g. Fig. 5 and Fig 6).
@@
The fact that barred galaxies tend to be more baryon-dominated has important consequences for the loci of barred and unbarred galaxies in the Tully-Fisher (TF) relation (see Fig.  17). Namely, barred galaxies are offset from unbarred galaxies in the TF relation (i.e. for the same $V_{\rm flat}$, barred galaxies have a larger $M_{\star}$ or $M_{\rm baryonic}$). Confirming this offset for barred galaxies in the observed TF relation would have important implications for the existence of dark matter, since such an offset cannot be readily explained in a modified gravity (e.g. MOND) framework.
@@ Appendix B Relation between different estimates of baryon-dominance
Figure 22: The average (solid lines) and standard deviation (shaded regions) of the ratio of stellar-to-dark matter mass within the half-mass radius as a function of time for the barred galaxies (excluding the interacting galaxies), in red, and for the unbarred galaxies, in blue. On average, for barred galaxies, the stellar mass dominates over the dark matter mass within the half-mass radius by $z\sim 0$, in contrast to unbarred galaxies which tend to be more dark matter dominated over all cosmic times.
[…]
@@ Appendix C The Baryonic Tully Fisher Relation
Figure 23: The Baryonic Tully Fisher relation for galaxies at $z=0$ in Auriga. Left: The grey dot-dashed line shows the relation given by fitting all galaxies together, while the shaded region indicates the intrinsic scatter $\sigma_{y}$. Right: The red (blue) dashed line shows the relation when fitting only the barred (unbarred) galaxies. The shaded regions indicate the uncertainties from boostrapping (see text). The scatter is slightly reduced for the BTFR as compared to the STFR and the barred and unbarred BTFR relations are less separated.